OpenStudy (anonymous):

suppose that f(x) = 4x+7 and g(x) =2x. what is a rule for g(f(x)? what is a rule for f(g(x))? suppose that f(x) = 4x+7 and g(x) =2x. what is a rule for g(f(x)? what is a rule for f(g(x))?

OpenStudy (aakashsudhakar):

I'm not sure what you mean by "What is a Rule", so I'll try to simply solve these if that's fine. When f(x) = 4x + 7 and g(x) = 2x, what is g(f(x)) and what is f(g(x))? g(f(x)) = g(4x + 7) = 2(4x + 7) = 8x + 14 f(g(x)) = f(2x) = 4(2x) + 7 = 8x + 7 Hope that helps!